For the hydrolysis of a salt of a weak acid and a weak base, show that
Kh=Kw/(Ka)(Kb)
1
Expert's answer
2020-02-20T07:44:18-0500

Let we have "MA" a salt of weak acid and weak base. When "MA" is dissolved in water, the interaction occurs between"M^+, A^-"  ion and water molecules.

"M^+ + A^- + H_2O \u2194 HA + MOH"

At equilibrium,"K_{eq} = \frac{[HA][MOH]}{[M^+][A^-][H_2O]}"

"\implies K_{eq} [H_2O] = K_h =\frac{ [HA][MOH]}{ [M^+][A^-] }" (where "K_h" is hydrolysis constant.)....(1)

Ionic product of water "K_w = [H^+] [OH^-]"

And ionization constant of weak acid "K_a =\frac{[H^+] [A^-]}{[HA]}"

Ionization constant of weak base "K_b =\frac{ [M^+] [OH^-]}{[MOH]}"

"\frac{K_w}{K_a \times K_b } = \frac{ [H^+] [OH^-] [HA] [MOH]}{ [H^+] [A^-] [M^+] [OH^-] 
}" "= \frac{[HA] [MOH]}{[A^-] [M^+]}" "= K_h" ..(from (1))

"K_h = \frac{K_w}{ K_a \times K_b}"